OpenStudy (anonymous):

What is the measure of one interior angle in a regular hexagon? 135° 120° 360° 720°

OpenStudy (anonymous):

@mathstudent55 @mathmath333

OpenStudy (anonymous):

@randy.jackson

OpenStudy (anonymous):

@RyGuy

OpenStudy (anonymous):

@babyblue123

OpenStudy (anonymous):

sum of exterior angles=360 each exterior angle of regular hexagon=360/6=60 each interior angle=180-60=?

OpenStudy (anonymous):

@Breezyb123

OpenStudy (anonymous):

@baking4lyf

OpenStudy (anonymous):

so the answer is 120

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

thank you i have more

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

What is m∠WVX?  33° 140° 100° 107°